OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

1.(y-y1)=m(x-x1) y+2=2(x-3) 2x-y-8=0 2. now 2x-y=8 2x/8-y/8=1 x/4-y/8=1 3.now take l.c.m 2x-y/8=1 2x-y=8 2x-y-8=0

OpenStudy (anonymous):

part 1 point slope formula is given by y-y1=m(x-x1) therefore y+2=2(x-3) y +2 = 2x - 6 y = 2x -6 -2 y = 2x - 8

OpenStudy (anonymous):

Thank you both for helping understand it. I got so confused with the question :)